# Long-Range Graph Wavelet Networks

### v1

@@ 6 Discussion and Conclusion
Empirically, LR-GWN achieves state-of-the-art performance on long-range benchmarks and remains competitive on short-range tasks. More importantly, LR-GWN establishes wavelet-based filters as interpretable and theoretically grounded tools for graph learning, laying the foundation for adaptive and more expressive architectures. With LR-GWN, we aim to advance both the theoretical foundations and practical applications of wavelet-based long-range graph representation learning.
@@ A.1 Graphs and Laplacians
For undirected graphs, $\mathcal{L}$} is symmetric and positive semidefinite, which guarantees an eigenvalue decomposition $\smash{{\mathcal{L}}={\bm{U}}{\bm{\Lambda}}{\bm{U}}^{\top}}$. Here $\smash{{\bm{U}}\in\mathbb{R}^{n\times n}}$ contains the eigenvectors as columns, forming an orthogonal basis for functions defined on the graph ($\smash{{\bm{U}}{\bm{U}}^{\top}={\bm{I}}}$), and $\smash{{\bm{\Lambda}}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})}$ collects the eigenvalues. The spectrum is bounded as $\smash{0=\lambda_{1}\leq\dots\leq\lambda_{n}\leq 2}$. Small eigenvalues correspond to smooth, slowly varying components, while large eigenvalues capture oscillatory, high-frequency behavior (Chung, 1997).
@@ A.2 Graph Fourier Transform
The graph Fourier transform (GFT) (Shuman et al., 2011; Sandryhaila & Moura, 2013) generalizes the classical discrete Fourier transform (DFT) (Bracewell, 2000) to signals supported on graphs. In the classical Euclidean setting, a signal is decomposed into complex exponentials, whose frequencies correspond to sinusoidal oscillations of different scales. On a graph, there is no notion of translation or frequency in the usual sense. Instead, the eigenvectors of the Laplacian play the role of Fourier modes, with the associated eigenvalues serving as the notion of frequencies.
@@ A.3 Wavelet Analysis
Wavelet theory extends Fourier analysis by enabling localized, multiresolution representations of signals (Mallat, 1989, 2009). Whereas Fourier bases are global and capture only frequency information, wavelets provide both spectral and spatial localization by combining scaling and translation operations.
@@ A.4 Wavelets on Graphs
Hammond et al. (2011) extend wavelet analysis to graphs by defining filters in the Laplacian spectral domain. Let $\smash{{\mathcal{L}}={\bm{U}}{\bm{\Lambda}}{\bm{U}}^{\top}}$ be the eigendecomposition of the (normalized) Laplacian, and let $\smash{\widehat{\psi}}$ and $\smash{\widehat{\phi}}$ denote a band-pass wavelet kernel and a low-pass scaling kernel, respectively (both applied elementwise to eigenvalues).
@@ A.5 Chebyshev Polynomial Approximations
A widely used choice is the Chebyshev polynomial basis (Hammond et al., 2011; Defferrard et al., 2016). For a filter $\smash{\widehat{g}}$, one approximates
@@ B.1 Propagation and Hop-Limits
The interpretation of polynomial filters as message-passing operators is immediate from the fact that ${\mathcal{L}}^{k}$ encodes walks of length $k$ (Balcilar et al., 2020). A degree-$\rho$ polynomial filter can therefore only aggregate information from nodes within $\rho$ hops of each target node. This locality is beneficial for efficiency, but also imposes a hard ceiling on the effective receptive field. Deep message-passing GNNs, which correspond to higher-degree polynomials, in principle extend this range but suffer from over-squashing (Alon & Yahav, 2021): as the neighborhood grows exponentially with hop count, long-range signals are compressed into fixed-dimensional node embeddings, leading to severe information loss. Thus, regardless of how coefficients are chosen, polynomial filters cannot escape their fundamentally local nature.
@@ B.2 Function Approximation and Smoothness
Theorem 1 (Markov’s Inequality (Markov, 1889)).
@@ B.3 Convergence to Discontinuous Filters
Theorem 2 (Slow Polynomial Convergence (Geisler et al., 2024)).
@@ Appendix C Related Work
Wavelets on Graphs. The Spectral Graph Wavelet Transform (SGWT) by Hammond et al. (2011) introduced cubic spline wavelets and an exponential scaling function at fixed scales, ensuring admissibility but lacking learnable parameters. SGWT approximates filters using finite-order Chebyshev polynomials. Graph Wavelet Neural Networks (GWNN) (Xu et al., 2019) extend SGWT with fixed exponential wavelets, i.e., $g(s\lambda_{i})=e^{s\lambda_{i}}$. While building on wavelet theory, this construction does not enforce the admissibility condition, since $g(0)=1$. DEFT (Bastos et al., 2023) follows SGWT and parametrizes wavelet filters using free-parameters Chebyshev polynomials via a combination of GNNs and MLPs. The coefficients, however, are not bound and, by construction, do not enforce the admissibility condition. ASWT-SGNN (Liu et al., 2024b) approximates wavelet filter coefficients using Jackson–Chebyshev polynomials. Similarly to the discussion above about GWNN and DEFT, the coefficients are not bound to satisfy $g(0)=0$. AGT (Cho et al., 2024) extends GWNN, and learns node-wise scales, either exactly or via approximation. As with the methods discussed so far, this approach does not enforce the admissibility condition, and thus does not guarantee that the resulting filters behave as band-pass wavelets. WaveGC (Liu et al., 2025) satisfies the admissibility condition by parametrizing the scaling function with odd-terms of a Chebyshev polynomial, and the wavelet function with even-terms. However, the separation of basis terms limits functional expressiveness, and the use of non-wavelet-interpretable message passing diminishes the benefit of using wavelets.
@@
Combining Spatial and Spectral Filters for Long-Range Interactions. Hybrid GNNs combine local spatial and global spectral information to enhance long-range modeling. Stachenfeld et al. (2020) propose a hybrid message-passing framework that sacrifices permutation equivariance, while Liao et al. (2018) incorporate spectral information via the Lanczos algorithm. More recently, Geisler et al. (2024) introduced a spatial-spectral parametrization to improve efficiency. Beyond the explicit combination of spatial and spectral filters, graph rewiring (Gasteiger et al., 2019; Arnaiz-Rodr´ıguez et al., 2022) enhance connectivity and mitigate over-squashing by modifying connectivity. Graph transformers (GTs) (Min et al., 2022; Dwivedi et al., 2021; Geisler et al., 2023) use global attention mechanisms to enable interactions between distant nodes. GTs often rely on spectral positional encodings paired with message passing (Rampášek et al., 2023; Geisler et al., 2023). Rampášek et al. (2023) introduce a framework for general, powerful and scalable graph transformers (GPS), which allow for the definition of a broad family of architectures that combine local message passing with global attention mechanisms. Similarly to these models, LR-GWN integrates local and global components. In our case, local aggregation is achieved through polynomial filters, while global propagation is handled via a spectral wavelet filter. However, the core methodology and motivation differ. LR-GWN is rooted in wavelet theory, with the aim of modeling low-pass and band-pass filters in a principled and interpretable manner. Graph transformers, by contrast, cover a broad family of architectures with varying attention mechanisms and are generally not designed with the same frequency-selective properties in mind.
@@ D.1 Filter Parametrization
where $T_{i}(\cdot)$ is the $i$-th Chebyshev polynomial. This formulation enables spatially localized filtering with computational efficiency. Alternative polynomial bases, such as Bernstein (He et al., 2021) or Jacobi polynomials (Wang & Zhang, 2022), could also be employed.
@@
where ${\bm{W}}_{\theta^{(l)}}\in\mathbb{R}^{z\times d^{(l)}}$ are learnable weights, and $d^{(l)}$ is the output dimensionality of layer $l$. The basis $\operatorname{GaussianSmearing}(\lambda):[0,2]^{k}\to\mathbb{R}^{k\times z}$ projects eigenvalues onto $z$ Gaussian radial functions, introducing a frequency cutoff $\lambda_{\text{cut}}<\lambda_{\text{max}}$ (Schütt et al., 2017). We follow Geisler et al. (2024) and apply a frequency-domain windowing function to regularize the spectral response. Element-wise, the filter is defined as $\smash{S_{\bm{\theta}^{(l)}}({\bm{\Lambda}})_{u,v}:=S_{v}(\lambda_{u};\bm{\theta}^{(l)})}$. Alternative spectral parametrizations, such as SpecFormer (Bo et al., 2023), could be used depending on application-specific needs.
@@ D.1.1 Shared and Independent Filter Configurations
Wavelet-based graph filters rely on a set of scales $\{s_{j}\}_{j=1}^{J}$ to capture multi-resolution information. Instead of using fixed scales as in Hammond et al. (2011), we introduce a learnable interpolation scheme in the shared filter configuration (Section˜D.1.1). The scale bounds are defined as $\log s_{\min}=\log(t_{\text{L}}/\lambda_{\text{max}})$ and $\log s_{\max}=\log(t_{\text{U}}\lambda_{\text{LP}}/\lambda_{\text{max}})$, where $t_{\text{L}}$ and $t_{\text{U}}$ are learnable lower and upper limits, and $\lambda_{\text{LP}}$ emphasizes low-frequency components. Assuming the maximum eigenvalue of the graph to be $\lambda_{\max}=2$ as it is usually the case in real-world graphs, we have $\log s_{\min}=\log t_{\text{L}},\log s_{\max}=\log t_{\text{U}}\lambda_{\text{LP}}$. The intermediate scales are uniformly spaced in log-space,
@@ D.2.1 Initialization
The wavelet filter parameters are initialized using standard Xavier initialization (Glorot & Bengio, 2010) for neural network weights and by setting $\omega_{0}=1$ and $\omega_{i}=0$ for all $i=1,\dots,\rho$ in the Chebyshev polynomials. While the spectral and spatial components could theoretically be designed to match a predefined wavelet, such as the Mexican hat wavelet, we observed no significant improvements from this approach. Consequently, we adopt conventional parameter initialization without enforcing wavelet-based constraints in our experiments.

### v3

@@ 6 Discussion and Conclusion
Empirically, LR-GWN achieves state-of-the-art performance on long-range benchmarks and remains competitive on short-range tasks. More importantly, LR-GWN establishes wavelet-based filters as interpretable and theoretically grounded tools for graph learning, laying the foundation for adaptive and more expressive architectures. LR-GWN currently assumes static graphs and fixed truncation levels. Extending it to dynamic graphs or adapting $k$ during training remains an open avenue.
With LR-GWN, we aim to advance both the theoretical foundations and practical applications of wavelet-based long-range graph representation learning.
@@ A.1 Graphs and Laplacians
For undirected graphs, $\mathcal{L}$} is symmetric and positive semidefinite, which guarantees an eigenvalue decomposition $\smash{{\mathcal{L}}={\bm{U}}{\bm{\Lambda}}{\bm{U}}^{\top}}$. Here $\smash{{\bm{U}}\in\mathbb{R}^{n\times n}}$ contains the eigenvectors as columns, forming an orthogonal basis for functions defined on the graph ($\smash{{\bm{U}}{\bm{U}}^{\top}={\bm{I}}}$), and $\smash{{\bm{\Lambda}}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})}$ collects the eigenvalues. The spectrum is bounded as $\smash{0=\lambda_{1}\leq\dots\leq\lambda_{n}\leq 2}$. Small eigenvalues correspond to smooth, slowly varying components, while large eigenvalues capture oscillatory, high-frequency behavior [10].
@@ A.2 Graph Fourier Transform
The graph Fourier transform (GFT) [43, 39] generalizes the classical discrete Fourier transform (DFT) [7] to signals supported on graphs. In the classical Euclidean setting, a signal is decomposed into complex exponentials, whose frequencies correspond to sinusoidal oscillations of different scales. On a graph, there is no notion of translation or frequency in the usual sense. Instead, the eigenvectors of the Laplacian play the role of Fourier modes, with the associated eigenvalues serving as the notion of frequencies.
@@ A.3 Wavelet Analysis
Wavelet theory extends Fourier analysis by enabling localized, multiresolution representations of signals [33, 34]. Whereas Fourier bases are global and capture only frequency information, wavelets provide both spectral and spatial localization by combining scaling and translation operations.
@@ A.4 Wavelets on Graphs
Hammond et al. [22] extend wavelet analysis to graphs by defining filters in the Laplacian spectral domain. Let $\smash{{\mathcal{L}}={\bm{U}}{\bm{\Lambda}}{\bm{U}}^{\top}}$ be the eigendecomposition of the (normalized) Laplacian, and let $\smash{\widehat{\psi}}$ and $\smash{\widehat{\phi}}$ denote a band-pass wavelet kernel and a low-pass scaling kernel, respectively (both applied elementwise to eigenvalues).
@@ A.5 Chebyshev Polynomial Approximations
A widely used choice is the Chebyshev polynomial basis [22, 11]. For a filter $\smash{\widehat{g}}$, one approximates
@@ B.1 Propagation and Hop-Limits
The interpretation of polynomial filters as message-passing operators is immediate from the fact that ${\mathcal{L}}^{k}$ encodes walks of length $k$ [4]. A degree-$\rho$ polynomial filter can therefore only aggregate information from nodes within $\rho$ hops of each target node. This locality is beneficial for efficiency, but also imposes a hard ceiling on the effective receptive field. Deep message-passing GNNs, which correspond to higher-degree polynomials, in principle extend this range but suffer from over-squashing [1]: as the neighborhood grows exponentially with hop count, long-range signals are compressed into fixed-dimensional node embeddings, leading to severe information loss. Thus, regardless of how coefficients are chosen, polynomial filters cannot escape their fundamentally local nature.
@@ B.2 Function Approximation and Smoothness
Theorem 1 (Markov’s Inequality [35]).
@@ B.3 Convergence to Discontinuous Filters
Theorem 2 (Slow Polynomial Convergence [18]).
@@ Appendix C Related Work
Wavelets on Graphs. The Spectral Graph Wavelet Transform (SGWT) by Hammond et al. [22] introduced cubic spline wavelets and an exponential scaling function at fixed scales, ensuring admissibility but lacking learnable parameters. SGWT approximates filters using finite-order Chebyshev polynomials. Graph Wavelet Neural Networks (GWNN) [47] extend SGWT with fixed exponential wavelets, i.e., $g(s\lambda_{i})=e^{s\lambda_{i}}$. While building on wavelet theory, this construction does not enforce the admissibility condition, since $g(0)=1$. DEFT [5] follows SGWT and parametrizes wavelet filters using free-parameters Chebyshev polynomials via a combination of GNNs and MLPs. The coefficients, however, are not bound and, by construction, do not enforce the admissibility condition. ASWT-SGNN [30] approximates wavelet filter coefficients using Jackson–Chebyshev polynomials. Similarly to the discussion above about GWNN and DEFT, the coefficients are not bound to satisfy $g(0)=0$. AGT [9] extends GWNN, and learns node-wise scales, either exactly or via approximation. As with the methods discussed so far, this approach does not enforce the admissibility condition, and thus does not guarantee that the resulting filters behave as band-pass wavelets. WaveGC [29] satisfies the admissibility condition by parametrizing the scaling function with odd-terms of a Chebyshev polynomial, and the wavelet function with even-terms. However, the separation of basis terms limits functional expressiveness, and the use of non-wavelet-interpretable message passing diminishes the benefit of using wavelets.
@@
Combining Spatial and Spectral Filters for Long-Range Interactions. Hybrid GNNs combine local spatial and global spectral information to enhance long-range modeling. Stachenfeld et al. [44] propose a hybrid message-passing framework that sacrifices permutation equivariance, while Liao et al. [27] incorporate spectral information via the Lanczos algorithm. More recently, Geisler et al. [18] introduced a spatial-spectral parametrization to improve efficiency. Beyond the explicit combination of spatial and spectral filters, graph rewiring [16, 3] enhance connectivity and mitigate over-squashing by modifying connectivity. Graph transformers (GTs) [37, 13, 17] use global attention mechanisms to enable interactions between distant nodes. GTs often rely on spectral positional encodings paired with message passing [38, 17]. Rampášek et al. [38] introduce a framework for general, powerful and scalable graph transformers (GPS), which allow for the definition of a broad family of architectures that combine local message passing with global attention mechanisms. Similarly to these models, LR-GWN integrates local and global components. In our case, local aggregation is achieved through polynomial filters, while global propagation is handled via a spectral wavelet filter. However, the core methodology and motivation differ. LR-GWN is rooted in wavelet theory, with the aim of modeling low-pass and band-pass filters in a principled and interpretable manner. Graph transformers, by contrast, cover a broad family of architectures with varying attention mechanisms and are generally not designed with the same frequency-selective properties in mind.
Our mention of graph transformers is intended purely for conceptual contrast; they were not included in experimental comparisons, as our focus is restricted to wavelet-based operators.
@@ D.1 Filter Parametrization
where $T_{i}(\cdot)$ is the $i$-th Chebyshev polynomial. This formulation enables spatially localized filtering with computational efficiency. Alternative polynomial bases, such as Bernstein [23] or Jacobi polynomials [46], could also be employed.
@@
where ${\bm{W}}_{\theta^{(l)}}\in\mathbb{R}^{z\times d^{(l)}}$ are learnable weights, and $d^{(l)}$ is the output dimensionality of layer $l$. The basis $\operatorname{GaussianSmearing}(\lambda):[0,2]^{k}\to\mathbb{R}^{k\times z}$ projects eigenvalues onto $z$ Gaussian radial functions, introducing a frequency cutoff $\lambda_{\text{cut}}<\lambda_{\text{max}}$ [41]. We follow Geisler et al. [18] and apply a frequency-domain windowing function to regularize the spectral response. Element-wise, the filter is defined as $\smash{S_{\bm{\theta}^{(l)}}({\bm{\Lambda}})_{u,v}:=S_{v}(\lambda_{u};\bm{\theta}^{(l)})}$. Alternative spectral parametrizations, such as SpecFormer [6], could be used depending on application-specific needs.
@@ D.1.1 Shared and Independent Filter Configurations
Wavelet-based graph filters rely on a set of scales $\{s_{j}\}_{j=1}^{J}$ to capture multi-resolution information. Instead of using fixed scales as in Hammond et al. [22], we introduce a learnable interpolation scheme in the shared filter configuration (Section˜D.1.1). The scale bounds are defined as $\log s_{\min}=\log(t_{\text{L}}/\lambda_{\text{max}})$ and $\log s_{\max}=\log(t_{\text{U}}\lambda_{\text{LP}}/\lambda_{\text{max}})$, where $t_{\text{L}}$ and $t_{\text{U}}$ are learnable lower and upper limits, and $\lambda_{\text{LP}}$ emphasizes low-frequency components. Assuming the maximum eigenvalue of the graph to be $\lambda_{\max}=2$ as it is usually the case in real-world graphs, we have $\log s_{\min}=\log t_{\text{L}},\log s_{\max}=\log t_{\text{U}}\lambda_{\text{LP}}$. The intermediate scales are uniformly spaced in log-space,
@@ D.2.1 Initialization
The wavelet filter parameters are initialized using standard Xavier initialization [20] for neural network weights and by setting $\omega_{0}=1$ and $\omega_{i}=0$ for all $i=1,\dots,\rho$ in the Chebyshev polynomials. While the spectral and spatial components could theoretically be designed to match a predefined wavelet, such as the Mexican hat wavelet, we observed no significant improvements from this approach. Consequently, we adopt conventional parameter initialization without enforcing wavelet-based constraints in our experiments.